# RTNinja: a generalized machine learning framework for analyzing random telegraph noise signals in nanoelectronic devices

### v2

@@ III.3 SourcesMapper
The SourcesMapper module is designed to establish the optimal $N$-source set that can explain the extracted 11 Bayesian levels along with the quantized signal (level-to-level transitions) information. An important assumption is made: each individual RTN source is considered to be independent and exhibits a two-state activity. This assumption implies that an $N$-source set can have a maximum of $2^{N}$ state configurations, representing superposition of the states of the sources. The objective is to establish the optimal $N$ sources, which includes determining the sources’ amplitudes (the difference between two distinct levels) and their activity, to achieve accurate reconstruction of the quantized data. A brute force approach would involve trying all combinations of possible amplitudes and finding the most likely fit for the sequence of transitions. This approach, however, becomes a combinatorial problem and is computationally infeasible to solve. Additionally, since a source’s amplitude is relative to the levels, it is necessary to determine the optimal baseline value, which is the signal that would be observed when all $N$ sources are in the off-state.
@@
The first step involves extracting meaningful features from the data generated by LevelsExtractor, as outlined in Figure LABEL:fig_transclust. The quantized data, representing level-to-level transitions, is modeled as a Markov process—a mathematical system in which transitions occur between states without memory of past states, a valid assumption for independent RTN signals. From this, a transition matrix is derived, containing the probabilities of transitions ($P_{T}$) between levels (Figure LABEL:fig_transmatrix). Simultaneously, the potential source amplitudes ($\Delta$) are calculated as the differences between the $\mu$ values of distinct levels. This results in a two-dimensional $P_{T}-\Delta$ space where unique groupings emerge, as illustrated in Supplementary Figure 1. These groupings reflect the distinct amplitude and activity characteristics of hidden sources. To extract an $N$-source set, the AP clustering algorithm is employed on a normalized $P_{T}-\Delta$ space to ensure equal-sized average error bars in both dimensions (Figure LABEL:fig_clustering). AP operates on the principle of ’message-passing’ between data points to identify distinct clusters [33]. It then automatically selects a data point from each cluster to act as the ’representative’ of that specific cluster. Unlike traditional clustering algorithms that require the number of clusters to be specified in advance, AP autonomously discovers both the clusters and their representatives. The results of the AP algorithm applied to the $P_{T}-\Delta$ space are shown in Figure LABEL:fig_clustering, where a total of 13 clusters are identified. Adjusting the AP hyperparameter settings led to slight variations in the number of clusters and representatives, as detailed in Supplementary Figure 1. Based on all these variants, an ensemble of one-dimensional representative potential amplitudes ($\Delta_{R}$) is prepared, without utilizing the representative $P_{T}$ information. To generate an $N$-source set, we first need to define the value of $N$. Given that the Bayesian levels are 11, the minimum number of hidden sources in the signal is 4 ($2^{4}=16$ state configurations). Multiple 4-source sets, referred to as candidate solutions, are then generated from the $\Delta_{R}$ ensemble, making the combinatorial number feasible ($\sim$700 in this case) for computational workload.
@@
(ii) The reconstruction accuracy with transition constraints metric assesses the validity of reconstructed quantized data under a specific rule: transitions between configurations must involve only the switching of a single source between on/off states. Simultaneous switching of multiple sources is considered a violation. This constraint is based on the fact that RTN sources can exhibit average activity rates ranging from the sampling rate to the full dataset length, for instance, from milliseconds to hours. Additionally, since RTN sources are considered independent, it is highly improbable for multiple sources to switch at the same instance. This approach helps maintain the fidelity of the reconstructed signal while accounting for realistic transition dynamics. Violations are counted during the reconstruction process to quantify the adherence to this constraint.
The two metrics are combined into a single normalized metric relative to a predefined tolerance limit, which is a hyperparameter. In this scenario, the tolerance limits are set such that up to 2% of total transitions are allowed as violations, and a mismatch tolerance of one standard deviation is permitted. Both metrics are equally weighted, resulting in a maximum cost function value of 2. An ideal candidate solution would perfectly match all Bayesian levels with no violations, yielding a cost function value of 0. Using this cost function, all candidate solutions are evaluated, and the optimal solution is identified as the one with minimum cost function value. In this case, the optimal solution has a cost function of 1.249. Figure LABEL:fig_sol illustrates the reconstructed RTN signal using this optimal candidate solution, which can be directly deconvolved into the activities of the four hidden sources (Figure LABEL:fig_solact). Since the dataset used in this example is synthetic and labeled, the RTNinja framework’s result is compared with the ground truth, demonstrating successful analysis of the convoluted RTN signal.
If no candidate solution achieves a cost function value below 2, the SourcesMapper module is automatically re-initialized under the hypothesis that the signal comprises one more source. This iterative process continues until the cost function condition is satisfied, ensuring an accurate representation of the RTN signal.
@@ IV Performance Evaluation
A comprehensive evaluation of the RTNinja framework’s performance across varying data complexities is presented in Figure 4(e) for three noise levels: 1%, 5%, and 30%. Additional results corresponding to intermediate noise levels of 10% and 20% are provided in Supplementary Figure 2. Figure LABEL:fig_confmatrix illustrates the comparison between the true number of sources in each dataset and the number estimated by RTNinja. The diagonal entries in this matrix indicate the number of datasets with correctly identified source counts. As noise levels increase, accurate detection becomes progressively more challenging due to weaker sources — i.e., those with low amplitudes — are increasingly suppressed by noise, often resulting in an underestimation of the true number of sources. Ideally, each row in the matrix should sum to 200 datasets, representing the complete batch for each source count. However, in practice, RTNinja occasionally fails to return a result for certain datasets (Supplementary Table 1). This is particularly true for datasets with fewer sources, where all sources become indistinguishable from noise due to low amplitudes. Conversely, in datasets with a higher number of sources and low noise levels, overfitting may occur due to the Bayesian algorithm overestimating the number of levels, which can lead to non-convergence in the SourcesMapper module. This is often caused by the initialization of the $N$-source hypothesis at an inflated value and the resulting computational burden of evaluating a large combinatorial space of candidate solutions. Hence, more datasets yield results at the highest noise level (30%) than at the lowest (1%) because the framework avoids overfitting in noisier conditions. Overall statistics on output yield across noise levels is provided in Supplementary Table 1.
@@
(e) Performance of the RTNinja framework at 1%, 5%, and 30% noise levels across all dataset complexities. (a) Matrix showing the number of datasets corresponding to each combination of true and estimated source counts, with diagonal entries indicating correct matches. (b) Activity match accuracy for mapped input-output sources, grouped by dataset complexity. The number of detected sources and probability of achieving >>95% activity match is indicated. (c) Accurate amplitude extraction is demonstrated, independent of noise level and data complexity. Blue squares indicate noise bounds; dashed lines mark $0.5\times$ and $2\times$ amplitude thresholds, with percentages showing sources within bounds. (d) Off-state and (e)On-state mean durations, with percentage of estimates within $0.5\times$-$2\times$ of true values. Accuracy is affected by inherent statistical sampling limitations and increasing noise, which can lead to underestimation.
@@
To rigorously assess RTNinja’s performance at the level of individual sources, we mapped each true (input) source to its corresponding estimated (output) source. This mapping is straightforward in cases where the number of estimated sources matches the true number of sources in the dataset. However, discrepancies arise when RTNinja underestimates or overestimates the number of sources — typically due to low-amplitude sources being missed in high-noise scenarios, or due to noise being mistakenly modeled as an additional source in low-noise, overfitted conditions. In such instances, we adopt an amplitude-based matching strategy: for example, if a dataset contains four true sources but RTNinja decomposes the signal into five sources, we select the four output sources whose amplitudes most closely match those of the true sources and discard the remaining unmatched component. This approach is similarly applied in cases of underestimation, where fewer sources are identified than are actually present. While the full temporal activity of each source is available, we deliberately rely solely on source amplitudes for input-output mapping to avoid artifacts that can arise from activity merging or splitting — a common side effect when the estimated number of sources deviates from the ground truth. This amplitude-centric matching ensures a consistent and interpretable evaluation even in the presence of estimation errors.
@@
Once the input-output source pairs have been successfully mapped, we assess RTNinja’s ability to accurately extract key characteristics of each individual source. The primary characteristic evaluated is the full activity profile of a source, measured using the activity match accuracy. To quantify this, we employ the Hamming distance, which calculates the dissimilarity between the binary activity patterns of the input and output, specifically whether a source is in the off- or on-state at each instance [34]. The Hamming distance is then converted into an activity match percentage by computing its complement. Figure LABEL:fig_gof illustrates the distribution of activity match percentages plotted against cumulative probability for each $N$-source dataset configuration across varying noise levels. Overall, RTNinja exhibits strong performance in recovering source activity. However, the accuracy of activity matching systematically decreases with increasing data complexity, and also declines with higher noise levels for any fixed number of sources. These trends highlight the increasing challenge of accurately capturing source-level dynamics under more complex and noisier signal conditions.
@@
Since the full binary activity trace of each source is available, we further analyze the durations of isolated on-state and off-state intervals. The temporal lengths of these continuous state segments approximate an exponential distribution, as discussed in Figure LABEL:fig_mcsim. We fit these distributions using a maximum likelihood estimation method to extract the primary parameters: the mean on-state duration ($\langle\tau_{off}^{n}\rangle$) and off-state duration ($\langle\tau_{on}^{n}\rangle$). The comparisons of these parameters between the ground truth and RTNinja outputs are shown in Figures LABEL:fig_toff and LABEL:fig_ton, respectively. RTNinja successfully recovers these distribution parameters for a substantial proportion of sources, as assessed by the percentage of mapped sources for which the extracted mean lies within a factor of two of the true value. However, accuracy declines systematically with increasing noise, accompanied by a growing tendency to underestimate the mean durations. This underestimation arises when some sources are not detected, causing their activity to be erroneously merged into the trajectories of detected sources — resulting in artificially faster dynamics. Additionally, the reliability of mean duration estimates is inherently limited by the number of observed transitions: for example, if a source enters the on-state only twice during the observation window, the estimated mean is constrained by statistical sampling limitations.
@@
Having evaluated RTNinja’s performance at the level of individual sources, we now step back to examine the framework’s aggregate performance across entire datasets, focusing specifically on how varying noise levels affect the results when data complexity is pooled. For each noise level, this involves analyzing the ensemble of all converged datasets (up to 1400 per condition, as detailed in Supplementary Table 1). Such an aggregated perspective is particularly relevant in the context of semiconductor device research, where the statistical characterization of RTN behavior is typically drawn from the overall distributions of defect parameters — most notably, amplitude and state duration metrics — and summarized using statistical descriptors like the mean and standard deviation [10]. These metrics are routinely employed to benchmark and compare device technologies in terms of their RTN impact.
@@
(f) Overall impact of noise level on RTNinja performance. (a) Histogram showing the frequency distribution of the difference between the estimated and actual number of sources across all datasets, grouped by noise level. A progressive leftward shift illustrates increasing underestimation as noise rises. (b) Cumulative probability of source activity match accuracy, based on amplitude-centric input-output mapping. The total number of mapped sources and the fraction with >>95% activity match is provided. (c) Complementary cumulative distribution function (CCDF) of estimated source amplitudes. Input (true) amplitude distributions are shown in grey as a reference. (d–e) Cumulative distributions of off-state and (e) on-state mean durations. True distributions are shown in grey. Deviations at higher noise levels reflect both missed sources and statistical compression due to limited sampling windows and temporal resolution.
@@
Figure LABEL:fig_ndiff presents the cumulative frequency of the difference between the estimated and actual number of sources across noise levels. As anticipated, a systematic shift toward underestimation is observed with increasing noise, reflecting a higher fraction of missed sources as the noise floor rises. This trend is further supported by the activity match results shown in Figure LABEL:fig_ngof, where activity accuracy — measured using the amplitude-based source matching strategy — declines inversely with noise. The primary source of this degradation lies in undetected sources whose activities are subsumed into the trajectories of detected ones. The overall distributions of mapped source characteristics — namely amplitudes, off-state mean durations, and on-state mean durations — are summarized in Figures LABEL:fig_nampdLABEL:fig_nton. RTNinja consistently recovers the underlying amplitude distributions with high accuracy, regardless of noise level. However, the total number of detected sources systematically declines as noise increases, particularly due to the suppression of low-amplitude components that fall below the elevated noise floor. At the other extreme, a minor deviation in amplitude estimation is occasionally observed for very high-amplitude sources. For both off-state ($\langle\tau_{off}^{n}\rangle$) and on-state ($\langle\tau_{on}^{n}\rangle$) mean durations, a clear trend of underestimation emerges with rising noise levels. This is primarily caused by the merging of undetected source activity into the traces of detected sources, leading to artificially faster dynamics. Additionally, both the lower and upper tails of the extracted duration distributions are truncated across all noise conditions — a limitation imposed by the finite resolution and observation window of the RTN signal.

### v3

@@ III.3 SourcesMapper
The SourcesMapper module is designed to establish the optimal $N$-source set that can explain the extracted 11 Bayesian levels along with the quantized signal (level-to-level transitions) information. An important assumption is made: each individual RTN source is considered to be independent and exhibits a two-state activity. This assumption implies that an $N$-source set can have a maximum of $2^{N}$ state configurations, representing the superposition of the states of the sources. The objective is to establish the optimal $N$ sources, which includes determining the sources’ amplitudes (the difference between two distinct levels) and their activity, to achieve accurate reconstruction of the quantized data. A brute-force approach would involve trying all combinations of possible amplitudes and finding the most likely fit for the sequence of transitions. This approach, however, becomes a combinatorial problem and is computationally infeasible to solve. In addition, since a source’s amplitude is relative to the levels, it is necessary to determine the optimal baseline value, which is the signal that would be observed when all $N$ sources are in the off-state.
@@
The first step involves extracting meaningful features from the data generated by LevelsExtractor, as outlined in Figure LABEL:fig_transclust. The quantized data, representing level-to-level transitions, is modeled as a Markov process—a mathematical system in which transitions occur between states without memory of past states, a valid assumption for independent RTN signals. From this, a transition matrix is derived, containing the probabilities of transitions ($P_{T}$) between levels [Figure LABEL:fig_transmatrix]. Simultaneously, the potential source amplitudes ($\Delta$) are calculated as the differences between the $\mu$ values of distinct levels. This results in a two-dimensional $P_{T}-\Delta$ space where unique groupings emerge, as illustrated in Supplementary Figure 1. These groupings reflect the distinct amplitude and activity characteristics of hidden sources. To extract an $N$-source set, the AP clustering algorithm is employed on a normalized $P_{T}-\Delta$ space to ensure equal-sized average error bars in both dimensions [Figure LABEL:fig_clustering]. AP operates on the principle of ”message-passing” between data points to identify distinct clusters [apc]. It then automatically selects a data point from each cluster to act as the ”representative” of that specific cluster. Unlike traditional clustering algorithms that require the number of clusters to be specified in advance, AP autonomously discovers both the clusters and their representatives. The results of the AP algorithm applied to the $P_{T}-\Delta$ space are shown in Figure LABEL:fig_clustering, where a total of 13 clusters are identified. Adjusting the AP hyperparameter settings led to slight variations in the number of clusters and representatives, as detailed in Supplementary Figure 1. Based on all these variants, an ensemble of one-dimensional representative potential amplitudes ($\Delta_{R}$) is prepared, without utilizing the representative $P_{T}$ information. To generate an $N$-source set, we first need to define the value of $N$. Given that the Bayesian levels are 11, the minimum number of hidden sources in the signal is 4 ($2^{4}=16$ state configurations). Multiple 4-source sets, referred to as candidate solutions, are then generated from the $\Delta_{R}$ ensemble, making the combinatorial number feasible ($\sim$700 in this case) for computational workload.
@@
(ii) The reconstruction accuracy with transition constraints metric assesses the validity of reconstructed quantized data under a specific rule: transitions between configurations must involve only the switching of a single source between on/off states. Simultaneous switching of multiple sources is considered a violation. This constraint is based on the fact that RTN sources can exhibit average activity rates ranging from the sampling rate to the full dataset length, for instance, from milliseconds to hours. In addition, since the RTN sources are considered independent, it is highly improbable for multiple sources to switch at the same instance. This approach helps maintain the fidelity of the reconstructed signal while accounting for realistic transition dynamics. Violations are counted during the reconstruction process to quantify adherence to the constraint.
The two metrics are combined into a single normalized metric relative to a predefined tolerance limit, which is a hyperparameter. In this scenario, the tolerance limits are set such that up to 2% of total transitions are allowed as violations, and a mismatch tolerance of one standard deviation is permitted. Both metrics are equally weighted, resulting in a maximum cost function value of 2. An ideal candidate solution would perfectly match all Bayesian levels with no violations, yielding a cost function value of 0. Using this cost function, all candidate solutions are evaluated, and the optimal solution is identified as the one with the minimum cost function value. In this case, the optimal solution has a cost function of 1.249. Figure LABEL:fig_sol illustrates the reconstructed RTN signal using this optimal candidate solution, which can be directly deconvolved into the activities of the four hidden sources [Figure LABEL:fig_solact]. Since the dataset used in this example is synthetic and labeled, the RTNinja framework’s result is compared with the ground truth, demonstrating successful analysis of the convoluted RTN signal.
If no candidate solution achieves a cost function value below 2, the SourcesMapper module is automatically reinitialized under the hypothesis that the signal comprises one more source. This iterative process continues until the cost function condition is satisfied, ensuring an accurate representation of the RTN signal.
@@ IV Performance Evaluation
A comprehensive evaluation of the RTNinja framework’s performance across varying data complexities is presented in Figure 4(e) for three noise levels: 1%, 5%, and 30%. Additional results corresponding to intermediate noise levels of 10% and 20% are provided in Supplementary Figure 2. Figure LABEL:fig_confmatrix illustrates the comparison between the true number of sources in each dataset and the number estimated by RTNinja. The diagonal entries in this matrix indicate the number of datasets with correctly identified source counts. As noise levels increase, accurate detection becomes progressively more challenging because weaker sources — i.e., those with low amplitudes — are increasingly suppressed by noise, often resulting in an underestimation of the true number of sources. Ideally, each row in the matrix should sum to 200 datasets, representing the complete batch for each source count. However, in practice, RTNinja occasionally fails to return a result for certain datasets (Supplementary Table 1). This is particularly true for datasets with fewer sources, where all sources become indistinguishable from noise due to low amplitudes. Conversely, in datasets with a higher number of sources and low noise levels, overfitting may occur due to the Bayesian algorithm overestimating the number of levels, which can lead to non-convergence in the SourcesMapper module. This is often caused by the initialization of the $N$-source hypothesis at an inflated value and the resulting computational burden of evaluating a large combinatorial space of candidate solutions. Hence, more datasets yield results at the highest noise level (30%) than at the lowest (1%) because the framework avoids overfitting in noisier conditions. Overall statistics on output yield across noise levels are provided in Supplementary Table 1.
@@
(e) Performance of the RTNinja framework at 1%, 5%, and 30% noise levels across all dataset complexities. (a) Matrix showing the number of datasets corresponding to each combination of true and estimated source counts, with diagonal entries indicating correct matches. (b) Activity match accuracy for mapped input–output sources, grouped by dataset complexity. The number of detected sources and probability of achieving >>95% activity match is indicated. (c) Accurate amplitude extraction is demonstrated, independent of noise level and data complexity. Blue squares indicate noise bounds; dashed lines mark $0.5\times$ and $2\times$ amplitude thresholds, with percentages showing sources within bounds. (d) Off-state and (e) on-state mean durations, with the percentage of estimates within $0.5\times$-$2\times$ of true values. Accuracy is affected by inherent statistical sampling limitations and increasing noise, which can lead to underestimation.
@@
To rigorously assess RTNinja’s performance at the level of individual sources, we mapped each true (input) source to its corresponding estimated (output) source. This mapping is straightforward in cases where the number of estimated sources matches the true number of sources in the dataset. However, discrepancies arise when RTNinja underestimates or overestimates the number of sources — typically due to low-amplitude sources being missed in high-noise scenarios or due to noise being mistakenly modeled as an additional source in low-noise, overfitted conditions. In such instances, we adopt an amplitude-based matching strategy: for example, if a dataset contains four true sources but RTNinja decomposes the signal into five sources, we select the four output sources whose amplitudes most closely match those of the true sources and discard the remaining unmatched component. This approach is similarly applied in cases of underestimation, where fewer sources are identified than are actually present. While the full temporal activity of each source is available, we deliberately rely solely on source amplitudes for input–output mapping to avoid artifacts that can arise from activity merging or splitting — a common side effect when the estimated number of sources deviates from the ground truth. This amplitude-centric matching ensures a consistent and interpretable evaluation even in the presence of estimation errors.
@@
Once the input–output source pairs have been successfully mapped, we assess RTNinja’s ability to accurately extract key characteristics of each individual source. The primary characteristic evaluated is the full activity profile of a source, measured using the activity match accuracy. To quantify this, we employ the Hamming distance, which calculates the dissimilarity between the binary activity patterns of the input and output, specifically whether a source is in the off- or on-state at each instance [hamming]. The Hamming distance is then converted into an activity match percentage by computing its complement. Figure LABEL:fig_gof illustrates the distribution of activity match percentages plotted against cumulative probability for each $N$-source dataset configuration across varying noise levels. Overall, RTNinja exhibits strong performance in recovering source activity. However, the accuracy of activity matching systematically decreases with increasing data complexity, and also declines with higher noise levels for any fixed number of sources. These trends highlight the increasing challenge of accurately capturing source-level dynamics under more complex and noisier signal conditions.
@@
Since the full binary activity trace of each source is available, we further analyze the durations of isolated on-state and off-state intervals. The temporal lengths of these continuous state segments approximate an exponential distribution, as discussed in Figure LABEL:fig_mcsim. We fit these distributions using a maximum likelihood estimation method to extract the primary parameters: the mean off-state duration ($\langle\tau_{off}^{n}\rangle$) and mean on-state duration ($\langle\tau_{on}^{n}\rangle$). The comparisons of these parameters between the ground truth and RTNinja outputs are shown in Figures LABEL:fig_toff and LABEL:fig_ton, respectively. RTNinja successfully recovers these distribution parameters for a substantial proportion of sources, as assessed by the percentage of mapped sources for which the extracted mean lies within a factor of two of the true value. However, accuracy declines systematically with increasing noise, accompanied by a growing tendency to underestimate the mean durations. This underestimation arises when some sources are not detected, causing their activity to be erroneously merged into the trajectories of detected sources — resulting in artificially faster dynamics. In addition, the reliability of mean duration estimates is inherently limited by the number of observed transitions: for example, if a source enters the on-state only twice during the observation window, the estimated mean is constrained by statistical sampling limitations.
@@
Having evaluated RTNinja’s performance at the level of individual sources, we now step back to examine the framework’s aggregate performance across entire datasets, focusing specifically on how varying noise levels affect the results when data complexity is pooled. For each noise level, this involves analyzing the ensemble of all converged datasets (up to 1400 per condition, as detailed in Supplementary Table 1). Such an aggregated perspective is particularly relevant in the context of semiconductor device research, where the statistical characterization of RTN behavior is typically drawn from the overall distributions of defect parameters — most notably, amplitude and state duration metrics — and summarized using statistical descriptors like the mean and standard deviation [grasser2020noise]. These metrics are routinely employed to benchmark and compare device technologies in terms of their RTN impact.
@@
(f) Overall impact of noise level on RTNinja performance. (a) Histogram showing the frequency distribution of the difference between the estimated and actual number of sources across all datasets, grouped by noise level. A progressive leftward shift illustrates increasing underestimation as noise rises. (b) Cumulative probability of source activity match accuracy, based on amplitude-centric input-output mapping. The total number of mapped sources and the fraction with >>95% activity match is provided. (c) Complementary cumulative distribution function (CCDF) of estimated source amplitudes. Input (true) amplitude distributions are shown in gray as a reference. (d–e) Cumulative distributions of off-state and (e) on-state mean durations. True distributions are shown in gray. Deviations at higher noise levels reflect both missed sources and statistical compression due to limited sampling windows and temporal resolution.
@@
Figure LABEL:fig_ndiff presents the cumulative frequency of the difference between the estimated and actual number of sources across noise levels. As anticipated, a systematic shift toward underestimation is observed with increasing noise, reflecting a higher fraction of missed sources as the noise floor rises. This trend is further supported by the activity match results shown in Figure LABEL:fig_ngof, where activity accuracy — measured using the amplitude-based source matching strategy — declines inversely with noise. The primary source of this degradation lies in undetected sources whose activities are subsumed into the trajectories of detected ones. The overall distributions of mapped source characteristics — namely, amplitudes, off-state mean durations, and on-state mean durations — are summarized in Figures LABEL:fig_nampdLABEL:fig_nton. RTNinja consistently recovers the underlying amplitude distributions with high accuracy, regardless of noise level. However, the total number of detected sources systematically declines as noise increases, particularly due to the suppression of low-amplitude components that fall below the elevated noise floor. At the other extreme, a minor deviation in amplitude estimation is occasionally observed for very high-amplitude sources. For both off-state ($\langle\tau_{off}^{n}\rangle$) and on-state ($\langle\tau_{on}^{n}\rangle$) mean durations, a clear trend of underestimation emerges with rising noise levels. This is primarily caused by the merging of undetected source activity into the traces of detected sources, leading to artificially faster dynamics. In addition, both the lower and upper tails of the extracted duration distributions are truncated across all noise conditions — a limitation imposed by the finite resolution and observation window of the RTN signal.